Hello,
Having a problem with
5/4x + 1/8x =3/8+x fractions 5 over 4 x + 1 0ver 8x = 3 over 8x
not sure how to do this, fractions are hard for me.
Laurie

Answer by rfer(16322)

5/4x+1/8x=3/8+x
10/8x+1/8x=3/8+8/8x
3/8x=3/8
x=(3/8)/(3/8)
x=3/8*8/3
x=24/24
x=1